# and even simpler proofs of quantumness

### v1

@@ 10.1 Compiler Guarantees
Let $\mathsf{G}$ be any contextuality game with ${\rm valNC}<1$. Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:C-1-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
@@
    ${\rm valQu}-\mathsf{negl}(\lambda).$
@@
    $1-\mathsf{const}_{1}+\mathsf{negl}^{\prime}(\lambda)\,,$
    where $\mathsf{const}_{1}=\min_{C\in C^{{\rm all}}}\Pr(C)/|C|=O(1)$, assuming $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad).
Completeness is straightforward to verify. We prove soundness in Section 11.
@@ 12.1 Compiler Guarantees
Let $\mathsf{G}$ be any contextuality game with ${\rm valNC}<1$ where all contexts are of the same size (i.e. $|C|=|C^{\prime}|$ for all $C,C^{\prime}\in C^{{\rm all}}$). Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:The-Cminus11-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
@@
    $1-\frac{1}{|C|}+\frac{{\rm valQu}}{|C|}-\mathsf{negl}(\lambda).$
@@
    $1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\mathsf{negl}^{\prime}(\lambda)\,,$
    assuming $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad).
Completeness is straightforward to verify. We prove soundness in Section 13.

### v2

@@ 11.1 Compiler Guarantees
Suppose $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad). Let $\mathsf{G}$ be any contextuality game with ${\rm valNC}<1$. Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:C-1-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
@@
    $c(\lambda):={\rm valQu}-\mathsf{negl}(\lambda).$
@@
    $s(\lambda):=1-\mathsf{const}_{1}+\mathsf{negl}^{\prime}(\lambda)\,,$
    where $\mathsf{const}_{1}=\min_{C\in C^{{\rm all}}}\Pr(C)/|C|=O(1)$.
Furthermore, $G^{\prime}_{\lambda}$ is faithful to $\mathsf{G}$ (as in LABEL:Def:OperationalTestOfContextuality) with parameters $s(\lambda)$ and $c(\lambda)$.
Completeness is straightforward to verify. The proof of faithfulness is analogous to that of LABEL:Thm:CompiledGameIsSecure. We prove soundness in Section 12.
@@ 13.1 Compiler Guarantees
Suppose $\mathsf{QFHE}$ and $\mathsf{OPad}$ are secure (as in LABEL:Def:QFHEscheme,Oblivious-U-pad), and compatible (as in LABEL:Def:QFHEcomptableopad). Let $\mathsf{G}$ be any contextuality game with ${\rm valNC}<1$ where all contexts are of the same size (i.e. $|C|=|C^{\prime}|$ for all $C,C^{\prime}\in C^{{\rm all}}$). Let $\mathsf{G}^{\prime}_{\lambda}$ be the compiled game produced by LABEL:Alg:The-Cminus11-compiler on input $\mathsf{G}$ and a security parameter $\lambda$. Then, the following holds.
@@
    $c(\lambda):=1-\frac{1}{|C|}+\frac{{\rm valQu}}{|C|}-\mathsf{negl}(\lambda).$
@@
    $s(\lambda):=1-\frac{1}{|C|}+\frac{{\rm valNC}}{|C|}+\mathsf{negl}^{\prime}(\lambda)\,,$
Furthermore, $G^{\prime}_{\lambda}$ is faithful to $\mathsf{G}$ (as in LABEL:Def:OperationalTestOfContextuality) with parameters $s(\lambda)$ and $c(\lambda)$.
Completeness is straightforward to verify. The proof of faithfulness is analogous to that of LABEL:Thm:CompiledGameIsSecure. We prove soundness in Section 14.